OpenStudy (anonymous):

What is the next number in the pattern: 5, 7, 6, 8, 7, 9, 8, …? A. 7 B. 9 C. 10 D. 11

OpenStudy (anonymous):

I think it's C

OpenStudy (hunus):

It's C

OpenStudy (anonymous):

5 + 2= 7 - 1= 6 + 2 = 8-1 = 7 ......

OpenStudy (anonymous):

Wait so is it 7 or 10? I'm confused...

OpenStudy (anonymous):

I'm pretty sure it's seven look, first number= 5 Next number= 5+2 Next number= (5+2)-1 Next number= ((5+2)-1)+2 Next number= (((5+2)-1)+2)-1

OpenStudy (hunus):

10